Xiao Ming and Xiao Hong originally had a total of 100 yuan. Xiao Ming now earns another 50 yuan, while Xiao Hong spent 30 yuan. At this time, Xiao Ming's money is exactly three times that of Xiao Hong. So how much does Xiao Ming and Xiao Hong have?


1
Expert's answer
2021-08-09T11:13:20-0400

Suppose that Xiao Ming originally had "x" yuan. Then Xiao Hong originally had "100-x."


"x+50=3(100-x-30)"

"x+50=210-3x"

"4x=160"

"x=40"


"x+50=90"

"100-x-30=30"

Now Xiao Ming has 90 yuan and Xiao Hong has 30 yuan.